# Generative AI as a lab partner: a case study

### v1

@@ Abstract
Generative AI tools, including the popular ChatGPT, have made a clear mark on discourses related to future work and education practices. Previous research in science education has highlighted the potential for generative AI in various education-related areas, including generating valuable discussion material, solving physics problems, and acting as a tutor. However, little research has been done regarding the role of generative AI tools in laboratory work, an essential part of science education, and physics education specifically. Here we show various ways in which high school students use ChatGPT during a physics laboratory session and discuss the relevance of using generative AI tools to investigate acoustic levitation and the speed of sound in air. The findings show agreement with previous research regarding the importance of educating students about the capabilities and limitations of using generative AI. Contrasting fruitful and problematic interactions with ChatGPT during lab sessions with seven lab groups involving 19 high school students made it possible to identify that ChatGPT can be a helpful tool in the physics laboratory. However, the teacher plays a crucial role in identifying students’ needs and capabilities of understanding the potential and limitations of generative AI. As such, our findings show that generative AI tools may handle some questions and problems and thus demonstrate their potential to help distribute teachers’ workload more equitably during laboratory sessions. Finally, this study serves as an important point of discussion regarding the ways in which students need support and training to efficiently utilize generative AI to further their learning of physics.
@@ I Introduction
The introduction of modern generative artificial intelligence (GenAI) tools, spearheaded by the public release of GPT and the chatbot ChatGPT [1], have led to a broad debate on their impact on education and future workforce needs [2, 3, 4, 5]. In physics education, recent studies involving GenAI tools like ChatGPT focus on comprehension and its ability to generate valuable (reliable or interesting, but wrong) answers to physics problems [6, 7, 8, 9]. Here, a need is identified for physics instructors to incorporate artificial intelligence (AI) in the classroom. In doing so, it is crucial to understand how students comprehend these tools in various situations, which has started to be investigated within the broader context of science education [10, 11]. However, studying natural science involves much experimental work, and thus, there is also a need to explore how GenAI can be utilized in laboratory settings to improve learning outcomes.
[…]
How, then, are students in STEM education able to make use of GenAI? In a study where first- and second-year undergraduate students were tasked to evaluate the scientific quality of answers, where three were wrong and generated using ChatGPT, and one correct was generated by the researchers but in a way to simulate how ChatGPT would present the answer [6]. There, it was identified that the students’ ability to gauge the quality of the answer was closely linked to their prior knowledge of the underlying physics. In another study where the potential of using ChatGPT as a tutor was investigated [11], students’ preconceptions about the capabilities of GenAI tools influenced their trust in answers provided by ChatGPT. Further, they also identified that many of the students in the study blindly trusted answers provided by ChatGPT. As such, current research into students’ views on GenAI shows both promising avenues for implementation and raises relevant concerns about the use of such tools.
I.1 Aim and Research questions
With the identified need to further explore in more nuanced ways how GenAI tools could be used in varying educational contexts and the intricate relationship between students and GenAI tools, our study aims to address this gap by reporting on a qualitative study where students explore the potential future role of GenAI tools, like ChatGPT, in the context of the high school physics laboratory. Specifically, our work provides valuable insight into how students are making use of ChatGPT during a lab session with varying prior knowledge of the phenomenon they are studying. To address the identified research gap, we aim to answer the following research questions (RQs)
@@ I.1 Aim and Research questions
II Method
@@
II.1 Experimental setup and lab design
In the field of acoustic levitation, particles typically in the millimeter range are subjected to forces that cause them to levitate. These forces are due to the creation of an acoustic field of standing waves, where it is possible to trap objects in the sound intensity nodes [17]. The experimental setup (LeviLab, [18]) comprises two ultrasonic loudspeakers mounted vertically and opposite to one another on a caliper (Fig. 1).
@@ II.1 Experimental setup and lab design
[…]
Two loudspeakers emitting sound waves of the same frequency create a standing wave at specific distances apart. This phenomenon is caused by constructive interference and results in an acoustic field with regions of low and high acoustic radiation pressure, also known as sound intensity nodes and antinodes. The pressure differential between these regions gives rise to the acoustic radiation force on macroscopic objects situated within the acoustic field, which can be sufficiently strong to counterbalance the gravitational force.
In the current study, students had access to LeviLab, a DC power supply, millimeter-sized polystyrene particles, tweezers, and a lab manual. Additionally, the students were encouraged to also bring their physics textbook, formula book, and any calculator. The manual included a list of materials, including information on what LeviLab consisted of. Regarding the data collection process, the lab manual described the procedure as follows. Place the speakers some distance apart and place a polystyrene particle on the lower speaker. Then, gradually lower the upper speaker using the caliper until the particle is observed to levitate. Note the distance between the speakers and then continue lowering the speakers until you have observed the particle levitate approximately ten times. When conducting the experiment, also take note of the ambient temperature of the room.
The lab manual also had assignments related to both data analysis and some final discussion questions. However, the instructions related to the analysis had been intentionally designed to be somewhat vague, asking students to perform a linear regression and to calculate the speed of sound using a formula that includes wavelength and frequency without explicating the exact form of the equation. Additionally, students were also tasked to compare the speed of sound with another model where the speed of sound depends on temperature. The intentional vagueness made it so that all or most groups would have to consult ChatGPT to identify what equation to use since it was beyond their expected prior knowledge, regardless of educational level. Finally, the discussion questions involved asking the students to construct a figure that described how sound waves can be used to counteract gravity, potential sources of errors, the relevance of measuring temperature, as well as applications of acoustic levitation. The discussion questions were added to further increase the need to consult ChatGPT during the lab session, but students were considered to have successfully finished the lab if they finished the analysis but not all discussion questions.
II.2 Participants
@@
Participating students taking Physics 1 or 2 were sought at one high school, and students taking Physics 3 were sought at a university that offered the non-mandatory course Physics 3 to third-year high school students. The sampling of students was done by convenience sampling, where all willing students got to participate in the study. In total, seven groups participated: four with second-year students taking Physics 1, one with third-year students taking Physics 2, and two with third-year students taking Physics 3. All groups consisted of three students, except Group 7 (Table 1).
@@ II.5 Analysis and limitations
The second step of the analytical process included inductively identifying relevant themes and passages that were in line with the general research aim of investigating the use of GenAI tools. Here, it was relevant to make use of the Gantt charts, transcripts, and each group’s chat history with ChatGPT to conclude first what they had done when, how they discussed interactions, or lack thereof, of ChatGPT, as well as connect why they answered the way they did during the post-interview in light of their reasoning during the pre-interview and interactions with ChatGPT during the lab. Following the second step, parts three and four of the analysis included identifying the overall narrative tone and relevant themes and data to include in the findings sections. Here, tentative findings were discussed both among all authors of the papers as well as presented and discussed with colleagues external to the current study. Finally, after having reached a consensus among the authors of the overall story identified in the data, the final analytical step includes reporting and discussing the findings, as is presented in the findings, discussion, and conclusion section of this paper.
@@ III.1 Group interviews
III.2 Interactions with ChatGPT during the lab
@@
[…]
@@
Groups 4 and 6 both did well during the lab and showed a deeper understanding of how to interact with ChatGPT during the lab. Questions they posed also included some poorly formulated ones such as “which sl unit is wavelength written with?”. However, a meaningful difference was the amount of follow-up and additional probing the interactions included. Group 6 notably chose to double-check the answers given by ChatGPT by opening up additional chats and asking similar questions again but without having the context of the entire chat history. Both groups also tried to more deeply understand how acoustic levitation works by posing different questions to ChatGPT, then discussed the answers among them and posed a new question to see if their understanding aligned with the new answers provided by ChatGPT. Such an interaction can be seen in Group 4 towards the end of the lab session (50 minutes into the session) when they discussed how particles could be levitated using the experimental setup together with ChatGPT using the prompts “how can sound waves counteract gravity”, “what do you mean by the gravitational force is balanced?”, “simplify how sound waves counteract the gravitational force ?”, and “is it particles that have bounced on each other and in that way creates a force upward that takes out the gravitational force”. In between the prompts, Group 4 tried to identify if the answers they got agreed with their prior knowledge and results from the experiment. However, due to their lack of adequate knowledge about the underlying physics, they ultimately struggled with identifying if the explanation ChatGPT gave as an answer to their final prompt was correct. Further, the answer mentioned sound waves, vibration, and energy transfer via collisions. It ultimately reinforced the students’ notion that it is more about some force pushing the particle from below than a more complex phenomenon involving standing waves.
IV Discussion

### v2

@@ Abstract
Generative AI tools, including the popular ChatGPT, have had a significant impact on discourses about future work and educational practices. Previous research in science education has highlighted the potential of generative AI in various education-related areas, including generating valuable discussion material, solving physics problems, and acting as a tutor. However, little research has been done regarding the role of generative AI tools in laboratory work, an essential part of science education, and physics education specifically. Here we show various ways in which high school students use ChatGPT during a physics laboratory session and discuss the relevance of using generative AI tools to investigate acoustic levitation and the speed of sound in air. Additionally, employing variation theory as a theoretical lens in the analysis, we identify how generative AI can be used to further develop students’ problem-solving skills in the physics laboratory. However, although our study shows that generative AI tools may handle some relevant questions and problems during laboratory work, the teacher still plays a crucial role in identifying students’ needs and capabilities of understanding the potential and limitations of generative AI. Finally, this study serves as an important point of discussion regarding the ways in which students need support and training to efficiently utilize generative AI to further their learning of physics.
@@ I Introduction
The introduction of generative artificial intelligence (GenAI) tools, spearheaded by the public release of the chatbot ChatGPT [1], has led to a broad debate on their impact on education and future workforce needs [2, 3, 4, 5]. In physics education research (PER), recent studies involving GenAI tools have examined their ability to generate valuable (i.e., reliable or interesting, but sometimes incorrect) answers to physics problems [6, 7, 8, 9]. Although a need has been identified for physics instructors to incorporate artificial intelligence (AI) in the classroom, there is still little focus on how GenAI tools can be used in the physics laboratory, a crucial part of students’ education when learning physics.
In this study, we investigate the potential future role of GenAI tools, like ChatGPT, in the context of high school physics laboratory. We do this in part by analyzing students’ awareness during the lab by drawing on the variation theory of teaching and learning as a theoretical framework. Specifically, our work provides valuable insights into how students are utilizing the freely available version of ChatGPT, GPT-3.5 at the time of data collection, with varying prior knowledge of the phenomenon they are studying, as well as with little to no prior training on how to use GenAI tools in school. Our aim is to contribute to the discussion on the ways in which GenAI tools may impact future laboratory work and implications for teachers, both in planning and executing such lessons. We do this by answering the following research questions (RQs):
@@
The structure of the paper is as follows. In Sec. II, we provide a broader context for the study by conducting a literature review to identify prior research that can help inform the use of GenAI in the physics laboratory. Sec. III describes the theoretical framework of variation theory, which is a theory of teaching and learning that was used in the current study. By introducing a theory of teaching and learning, it provides a clear lens through which we can connect our results to practical implications for teaching and learning in the physics laboratory. In the current study, we draw upon our theoretical framework specifically in analyzing the results as a way to answer RQ3. Our lab design, as well as the study design, is presented in Sec. IV, which also includes a discussion of the study’s limitations and how the data were analyzed. Our findings are presented in Sec. V. In Sec. VI, the RQs are discussed in relation to our findings. There, we also discuss implications for teaching and learning in the science laboratory. Finally, Sec. VII provides a summary of the paper, including an outlook that relates the findings of the current paper to future research needs in the field.
II Background
PER involving modern GenAI tools (post public release of ChatGPT) has had a broad focus, including comparing different models’ performance on tasks (e.g. [10, 11]), performance on physics problems (e.g. [8, 12, 13, 9, 14, 15]), ways in which physics teachers can use these tools (e.g. [7, 16, 17, 18, 19, 20]), and the potential for interactions between students and GenAI (e.g. [6, 21, 22, 23, 24]). A common theme that can be observed in PER is that the use of GenAI tools, whether for students, teachers, or researchers, must be carefully considered to achieve a beneficial outcome. Regarding the potential use of GenAI tools for students, a primary benefit has been identified to be connected to increased motivation and engagement [25]. However, a recent systematic review on the influence of ChatGPT on student engagement has found that the increased engagement can, in some instances, be linked to cheating and reduced critical thinking, resulting from over-reliance on answers provided by chatbots during problem-solving activities [26]. These findings reinforce the notion that the use of such tools in educational contexts requires careful consideration.
In science education, research about incorporating GenAI tools in laboratory work has primarily been done in the context of chemistry education, including comparing different models [27], aiding students developing methods [28], producing code to assist in data analysis [29], and in writing lab reports [30]. Two studies, both conducted at the university level and utilizing ChatGPT with the freely available version, GPT-3.5, involved students who had little to no prior experience with GenAI tools [29, 28]. In using ChatGPT to help students produce code to analyze lab results, students found the GenAI tool to be useful and reported an increased level of understanding of the programming language they used [29]. The other study had students use ChatGPT to help them develop methods for lab proposals in analytical and inorganic chemistry [28]. Results indicated that it was challenging to obtain useful information from ChatGPT about specific procedures, as it often provided some correct information alongside vague or incorrect information. Additionally, the authors found that students with more theoretical and experimental knowledge were able to utilize ChatGPT more effectively than novice students.
A link between students’ prior knowledge and their ability to evaluate the quality of information provided by GenAI tools has also been identified in PER, outside laboratory contexts. In a study investigating students’ ability to evaluate the scientific quality of answers to three physics problems, 102 first- and second-year undergraduate students were presented with four different answers to each problem, three incorrect and generated using ChatGPT, and one correct generated by the researchers [6]. Their findings identified that the students’ ability to gauge the quality of the answer was closely linked to their prior knowledge of the underlying physics.
Finally, in addition to prior knowledge influencing students’ ability to identify incorrect information provided to them by GenAI tools, their implicit trust in chatbots can also influence how they gauge answers they get when interacting with the tool. This was investigated in a study having students use ChatGPT as a tutor [21]. They found that students preconceptions about the capabilities of GenAI tools influenced their trust in answers provided by ChatGPT. Additionally, they also identified that many of the participating students blindly trusted answers provided by ChatGPT, pointing towards students not having received much in terms of prior teaching about GenAI tools or how they are typically developed.
III Variation theory
The current study is exploratory in nature, and the findings presented in this paper were identified using an inductive approach, similar to when applying grounded theory [31] as a research approach, or conducting a thematic analysis [32]. In this section, we introduce the variation theory of teaching and learning as a theoretical framework to be used in discussing our findings, allowing for further analysis related to students’ awareness, as well as discussing general principles on teaching and learning that the theory identifies. As such, variation theory provides a lens through which we can discuss the implications of our findings for teaching and learning in the science laboratory.
Variation theory is a theory of teaching and learning asserting that for learning to take place, the learner must notice what is to be learned, and that noticing requires variation [33, 34, 35, 36]. It argues that it is easier to notice what differs between two contexts than what is similar, and learning is to be understood as ‘learning to see’ [35]. Here, we can identify that learning is closely connected to a learner, meaning that in variation theory, one must acknowledge that the process of learning about a specific concept or phenomenon within a particular subject can not be viewed as an isolated event, void from individual differences, including previous experiences of what is to be learned. Rather, a teacher must account for the students when designing teaching sequences, because students’ prior knowledge and experiences of the world affect how they discern the current content.
III.1 Object of learning, structure of awareness, and critical aspects
What is to be learned is referred to as an object of learning. It is related to a specific group of students and the surrounding educational context, including curriculum-imposed goals [35, 36]. An object of learning is constituted by a direct and an indirect object of learning. The content (e.g., standing waves) to be learned refers to the direct, whereas a particular skill or capability to be developed (e.g., predicting how particles can be trapped using acoustic levitation) refers to the indirect object of learning [35]. Together, they make up the whole object of learning, e.g., using the concept of standing waves to explain and predict how to trap particles using acoustic levitation, which is then identified as what is to be learned.
For a particular object of learning, one must also consider how a particular group of students may initially perceive or discern it. How a student perceives an object of learning is related to their current structure of awareness, an idea stemming from the phenomenographic tradition [37]. In the phenomenographic tradition, it assumes that a learner’s structure of awareness is their experiences of the world. The structure of awareness continuously changes, as humans can only be consciously aware of a limited set of the world simultaneously. When considering an object of learning, we are thus never capable of being actively aware of all parts, or aspects, of it at the same time. As a student, what is to be learned can, from this point of view, be understood as becoming aware of parts of the object of learning in more powerful ways. These parts, that the student has yet to discern, are known as critical aspects [37, 38], and the structure of awareness dictates how students become aware of these aspects [39]. Finally, the object of learning can be divided into an intended (the teacher’s aim prior to the current lesson), an enacted (what aspects are made possible to discern during the lesson), and a lived object of learning (what the students actually learned).
III.2 Putting it into practice
When planning teaching and learning situations, variation theory emphasizes that focus should be on how variation can be used in handling the content that is taught, not variation in terms of varying teaching methods, such as types of classroom activities [36]. By varying the handling of the content, one can highlight the critical aspects of the object of learning in ways that make them more clearly visible to the students. However, if the teacher is not cognizant of their students’ current possible structure of awareness, introduced variations may shift their focus towards unproductive, misleading aspects of the phenomenon under study. Specifically, during a lesson, students are ideally better at discerning critical aspects of the phenomenon under study towards the end of the lesson. This means that the students have become more proficient in identifying what the important (critical) aspects are. The students’ capacity to better connect the critical aspects to the object of learning means that they have developed a more refined awareness of the phenomenon [39]. If students become aware of the intended object of learning during the lesson, it also means that the enacted object of learning is in line with the intended. Such an alignment relies on the students discerning the critical aspects, but also that the students’ structure of awareness is directed towards the object of learning.
Studies employing variation theory commonly focus on analyzing qualitative data, and the same holds for studies using this theory in PER (e.g., [40, 41, 42]). The three studies mentioned here to exemplify the use of variation theory in PER employ similar setups in that they follow a study design where a lesson has been designed. However, [40, 41] both have one set intervention design, whereas [42] iterates the design throughout the study, following a design-based research approach to study students’ understanding and improve the lesson design. Results from all three studies provide fruitful insights into what learning is made possible with each specific lesson design, and how teachers’ intervention can aid in further directing students’ attention towards critical aspects. On a more general note, they also highlight how variation theory can serve as a powerful tool in designing and planning good teaching and learning sequences in physics.
In the context of the current study, variation theory allows us to identify and discuss the implications of the findings presented in this paper for designing fruitful teaching and learning sequences in the age of AI in education. This is done in part by identifying instances during the lab where students’ awareness shifts, indicating that their conception of the object of learning, or some critical aspect(s), has changed.
IV Method
@@
IV.1 Experimental setup and lab design
In acoustic levitation, particles typically in the millimeter range are trapped mid-air due to the creation of an acoustic field of standing waves. Millimeter-sized particles consisting of materials denser than air are attracted to the sound intensity nodes [43]. The experimental setup (LeviLab, [44]) comprises two ultrasonic speakers mounted vertically and opposite to one another on a caliper (Fig. 1).
@@ IV.1 Experimental setup and lab design
As the speakers, which are connected in phase, are moved axially, the standing waves between them shift. If a millimeter-sized object is placed resting on the bottom speaker grill and the top speaker is slowly moved closer, the particle will be attracted to the closest node, which can either be above or below the speaker grill. This results in that the particle will, at certain distances when a node above the speaker grill is sufficiently close, seem to jump up and begin to hover mid-air. The configuration allows for the variable and precise measurement of the distance between the loudspeakers. In addition, LeviLab outputs a constant frequency of $40\text{ kHz}$, and is equipped with a thermometer and a digital display, which allows a measurement of the ambient temperature while performing experiments using the setup.
Additionally, the students were encouraged to also bring their physics textbook, formula book, and any calculator. The lab manual guided the students in obtaining data to analyze, from which they could calculate the wavelength and the speed of sound. The method for data collection was precisely described. However, the instructions related to the analysis had been intentionally designed to be somewhat vague, asking students to perform a linear regression and to calculate the speed of sound using a formula that includes wavelength and frequency without explicating the exact form of the equation. Additionally, students were also tasked to compare the speed of sound with another model where the speed of sound depends on temperature. The intentional vagueness made it so that all or most groups would have to consult ChatGPT to identify what equation to use, since it was beyond their expected prior knowledge, regardless of educational level. Lastly, to increase the need to consult ChatGPT, discussion questions were added, asking the students to construct a figure that describes how sound waves can be used to counteract gravity, potential sources of errors, the relevance of measuring temperature, and regarding applications of acoustic levitation. Students were considered to have successfully finished the lab if they finished the analysis, even if not all discussion questions were answered.
IV.2 Participants
@@
Participating students taking Physics 1 or 2 were sought at one high school, and students taking Physics 3 were sought at a university that offered the non-mandatory Physics 3 course to third-year high school students. The sampling of students was done by convenience sampling, where all willing students got to participate in the study. In total, seven groups participated: four with second-year students taking Physics 1, one with third-year students taking Physics 2, and two with third-year students taking Physics 3. All groups consisted of three students, except Group 7 (Table 1).
@@ IV.5 Analysis and limitations
The second step of the analytical process included inductively identifying relevant themes and passages that were in line with the general research aim of investigating the use of GenAI tools. Here, it was relevant to make use of the Gantt charts, transcripts, and each group’s chat history with ChatGPT to conclude first what they had done when, how they discussed interactions, or lack thereof, of ChatGPT, as well as connect why they answered the way they did during the post-interview in light of their reasoning during the pre-interview and interactions with ChatGPT during the lab. Following the second step, parts three and four of the analysis included identifying the overall narrative tone and relevant themes and data to include in the findings sections. Here, tentative findings were discussed both among all authors of the papers as well as presented and discussed with colleagues external to the current study. Then, the main author proposed an initial structure of the findings and the main conclusions to be presented, which were again discussed among all authors to certify that the identified conclusions are supported by the data. Practically, this was achieved by proposing themes and supporting transcripts, which were then discussed in terms of their reliability and potential problematic subjective interpretations. After reaching a consensus among the authors on the main emerging themes identified in the data, the final analytical step involves reporting and discussing the findings, as presented in Sec. V through Sec. VII of this paper.
In discussing the findings, the theoretical framework of variation theory was used in further analysis, related to the students’ awareness. Specifically, we relied on the structure of awareness as a way of understanding, from the students’ point of view, the enacted object of learning. The setup of the study, where no teacher was present to enact the intended object of learning, made it relevant to conceptualize the enacted object of learning. In our analysis, we treated the enacted object of learning as something dynamic, akin to the dynamical nature of the structure of awareness. As such, we could identify differences in what the enacted object of learning was between the groups, which aided our understanding of why the groups had seemingly different foci during the lab. Additionally, variation theory was used as a lens in the discussion for identifying implications for teaching and learning.
@@ V.1 Group interviews
V.2 Interactions with ChatGPT during the lab
@@
In general, groups 1, 2, 3, 5, and 7 were identified as struggling during the lab, meaning that they had trouble understanding what data to collect, got stuck during the analysis, or both. For groups 1, 2, 3, and 5, interactions with ChatGPT focused on asking simple questions and they did little to no follow-up. Group 7 struggled mainly due to a lack of knowledge on how to use the online calculator Desmos. They opted for using Desmos since they did not bring a calculator to the lab session, and thus, much of the interaction with ChatGPT focused on trying to get help with understanding how to use Desmos to perform the necessary analysis.
@@
Groups 4 and 6 both did well during the lab and showed a deeper understanding of how to interact with ChatGPT during the lab. Questions they posed also included some poorly formulated ones such as “which sl unit is wavelength written with?”. However, a meaningful difference was the amount of follow-up and additional probing the interactions included. Group 6 notably chose to double-check the answers given by ChatGPT by opening up additional chats and asking similar questions again but without having the context of the entire chat history. Both groups also tried to more deeply understand how acoustic levitation works by posing different questions to ChatGPT, then discussed the answers among them and posed a new question to see if their understanding aligned with the new answers provided by ChatGPT.
V.3 On interacting with ChatGPT and students awareness
Drawing upon our theoretical framework, we can identify instances where the students’ interactions with ChatGPT during the lab impacted their structure of awareness, which in turn has an effect on what the object of learning is (and thus what aspects are relevant to consider). As an initial step, it must be acknowledged what students perceived to be the primary focus at the beginning of the lab session. An aim for the lab session was provided to students by including a short sentence in the lab manual (see Appendix A) stating that they should explore acoustic levitation and determine the speed of sound in air. During the data collection phase of the lab, both groups 3 and 6 verbally reiterated the written aim. This led to two different interactions with the goal of better understanding acoustic levitation, where Group 6 discussed among themselves and reasoned about the phenomenon of acoustic levitation based on their prior physics knowledge. Group 3 instead approached the aim of exploring acoustic levitation by having Carl interact with ChatGPT, as described above. For the other groups, no discussions touching on the aim of exploring acoustic levitation were present until either during the data analysis phase of the lab, or after having finished the analysis. Among these, Group 4 stood out by having conceptual discussions on why the particle could levitate during the data analysis stage. Overall, most students perceived the aim of the lab session to be related to determining the speed of sound, and thus, the enacted object of learning was identified to be more closely connected to the capability of determining the speed of sound using the available equipment. In light of the identified enacted object of learning, it was relevant for the groups to spend most of their time on data collection and analysis, as illustrated by Fig. 3.
From the previous subsection presenting how groups interacted with ChatGPT, it became evident that the interactions were akin to having quick access to the teacher, a teaching assistant, or web searching capabilities. In these interactions, we can identify passages where students’ awareness of what could be important to discern during the lab was seemingly affected. Here we present three examples, from groups 2, 4, and 5, highlighting how the interactions affected the students’ structure of awareness during the episodes.
The first episode is of Group 2 and an initial interaction with ChatGPT during the data analysis phase. They asked ChatGPT “How to convert kelvin to celcius”, with the follow-up “how many degrees celdius is 273 kelvin”. In the first response, ChatGPT provided the formula $T_{C}=T_{K}-273.15$, and also included an example where $T_{K}=300\text{ K}$. In the second response, ChatGPT said that with $T_{K}=273\text{ K}$, $T_{C}=-0.15\text{ \textdegree C}$. After the replies, the group went on to have a discussion without ChatGPT about whether they needed the temperature data or not for the experiment. Here, the group still had no equation or conception of a model that included the temperature dependency for the speed of sound, making the interaction episode rather irrelevant. The students’ focus was locked on the aforementioned enacted object of learning, and they had yet to succeed with the first part of the analysis, which was to calculate the wavelength. However, with no apparent understanding of how the speed of sound, or how the concept of wavelengths were connected to the experimental setup, Group 2 would have needed clearer guidance towards aspects relevant to the object of learning. Instead, here the group’s interaction with ChatGPT strengthened their focus on temperature, and they kept on haphazardly trying various calculations to finish the lab.
In episode two, Group 5 also struggled with the analysis and getting an estimate of the speed of sound using the collected data. Towards the end of the lab session, they turned to ChatGPT and asked “what is speed of sound in 22 degrees celsius”. The reply told them that the approximate speed of sound at 22 $2\text{ \textdegree C}$ is $344\text{ m/s}$, and the group was then happy to not continue with their own calculations but simply agreed that the answer given by ChatGPT was a good estimation of what they would have ended up with. During the group’s short discussion following this interaction, they said that the reply was reasonable since ChatGPT provides “a good average of previous research results” (Eddie, Group 5), and the answer could thus be understood as ChatGPT had done the analysis for them. Here we could identify a similar enacted object of learning, albeit more closely connected to simply being able to provide an acceptable value for the speed of sound in air. Similar to Group 2, Group 5 showed no prior understanding of the speed of sound or the connection to the concept of wavelength. Drawing upon the identified enacted object of learning, together with the fact that Group 5 saw tools like ChatGPT as being able to provide a synthesis of all available information, their way of interacting with ChatGPT was perfectly reasonable. Again, here we see yet another episode where the students’ structure of awareness, both regarding the object of learning and GenAI tools, is not affected by spontaneous interactions with ChatGPT.
Episode three consists of an interaction Group 4 had with ChatGPT towards the end of the lab session. There, they discussed how particles could be levitated using the experimental setup together with ChatGPT using the prompts “how can sound waves counteract gravity”, “what do you mean by the gravitational force is balanced?”, “simplify how sound waves counteract the gravitational force ?”, and “is it particles that have bounced on each other and in that way creates a force upward that takes out the gravitational force”. In between the prompts, Group 4 tried to identify if the answers they got agreed with their prior knowledge and results from the experiment. For Group 4, we identified a clear contrast to the previous two episodes, namely that the enacted object of learning is more closely connected to the full stated aim in the lab manual. This is evident in part due to the fact that the group thought it important to not only calculate the speed of sound, but also to understand connections to the phenomenon of acoustic levitation. Another contrast is seen in the way Group 4 interacted with ChatGPT. Here, the group’s discussions prior to and between the first two interactions indicate that they initially have discerned a connection between the concept of wavelength and the levitating styrofoam particles. However, the students perceive that ChatGPT provides confounding information and introduces concepts and ideas they are currently unfamiliar with. Additionally, their lack of adequate knowledge about the underlying physics makes it hard for them to gauge the correctness of the answers ChatGPT provides. Thus, the interaction led to a shift of awareness regarding critical aspects of acoustic levitation with the equipment; from having to do with sound waves from the two speakers sometimes “strengthening” each other (I.e., on their way to discovering standing waves) and towards one students original notion that the styrofoam levitates mainly due to “air particles” physically pushing on it from below.
VI Discussion